Find the present value if a cash flow of $400 is received at the end of each quarter over the next 5 years with a discount rate of 8% p.a.
Answer #1

PV of cash flows of 400
Rate per quarter = 8%/4 = 2%
Number of Periods = 4*5 = 20
PV of Cash Flows using Annuity formula = PMT*(1-(1+r)-n)/r = 400*(1-(1+2%)-20)/2% = 6540.57

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well